# A poset game on submonoids of additively indecomposable ordinals

### v2

@@ Abstract.
Inspired by García-Marco and Knauer [GMK18], we consider the Chomp game on a natural well-partial order structure defined on submonoids $\mathcal{S}^{\sigma}\subseteq\omega^{\sigma}$ of additively indecomposable ordinals generated by sets of ordinals under natural sum and product of ordinals. We have called the generatinng sets of ordinals for which the game terminates in a finite number of steps well-partial order generators, and, under very mild conditions, we have characterized those that are finite.
A fundamental observation is that, for every ordinal $\sigma$, there exists an ordinal $\xi$ such that, for any class $\{(\mathcal{S}^{\tau};\leq_{\mathcal{S}^{\tau}})\mid\tau\geq\sigma\}$ of well-partial orders generated by a subset of $\omega^{\sigma}$, the second player has a winning strategy on all those posets or only on $\{(\mathcal{S}^{\beta};\leq_{\mathcal{S}^{\beta}})\mid\beta\in\alpha\}$ for some successor ordinal $\alpha\leq\xi$ (and the first player will have a winning strategy on the rest of the posets). This Hanf number-style property could be valuable in proving the existence of winning strategies. We conjecture that $\xi$ is the smallest possible and, using results from [Bag12], we prove that, if we restrict ourselves to the classes generated by finite sets of ordinals, $\xi<|\omega^{\sigma}|^{+}$. We also explicitly describe a winning strategy for a specific family of classes.
@@ 1. Introduction
In a poset game, two players take turns choosing elements from a partially ordered set and, on each turn, removing the chosen element and all those above it. The first player to run out of elements to choose loses. A particular case (and one of the most popular poset games) is the game Chomp: the poset is a rectangle of dimension $n\times m$ where each element is smaller than all the elements above it and to its right; the first player to eliminate the minimum (the element at the bottom left) loses. We can also define this game on any poset with a minimum and, of course, by removing this minimum, we obtain a poset game with the rules we explained at the beginning. The history of this combinatorial game dates back to 1952, when the Dutch mathematician Frederik Schuh introduced the “game of divisors”. Quite some time later, David Gale described in [Gal74] the Chomp game in the way just mentioned.
@@
Several generalizations of the Chomp game have been studied since its invention. In [GMK18], the Chomp game is studied on the posets given by numerical semigroups $\mathcal{S}\subset\mathbb{N}$ with the order defined by $a\leq_{\mathcal{S}}b$ iff $b-a\in\mathcal{S}$. This is a huge generalization of the original game, so that, as in the classic Chomp, there is no known constructive way to find winning strategies, except in particular cases.
@@
In [HS02], the original Chomp game is extended by allowing rectangles of the form $\alpha_{1}\times\cdots\times\alpha_{d}$, where $\alpha_{1},\dots,\alpha_{d}$ are ordinals. In the same spirit, here we propose a generalization of the game studied in [GMK18] to a set-theoretic context: as in $\mathbb{N}$, any additively indecomposable ordinal has a monoid structure on which we can define a partial order and, thus, play the Chomp game. In particular, we will focus on submonoids with natural sum (also called Hessenberg sum). The following sections will make clear that this family of submonoids is very interesting for our purposes. For example, we prove that it is possible to apply results from [GMK18] to this context to obtain constructive winning strategies. Apart from that, we will see that extending the problem to this context allows us to use purely set-theoretic techniques to draw some conclusions about the game.
@@
Since this paper may attract readers interested in game theory or set theory, the following section presents some very basic facts from both disciplines that will be necessary later. In the third section, we study some order-theoretic properties of the well-partial order generators, that is, sets of ordinals that generate monoids in which we can define a natural well-partial order structure. In particular, we provide a very general test to determine whether a finite set of ordinals is a well-partial order generator. In the final section, we will investigate classes of well-partial orders $\{(\mathcal{S}^{\tau};\leq_{\mathcal{S}^{\tau}})\mid\tau\geq\sigma\}$ generated by sets of ordinals $\Gamma$, where the sets $\mathcal{S}^{\tau}\subseteq\omega^{\tau}$ are monoids under natural sum. In particular, if the second player does not have a winning strategy on all the posets of the corresponding class, then there exists a successor ordinal $\alpha+1$, which we denote by $\text{ch}(\Gamma)$, such that the second player has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$ but not on $(\mathcal{S}^{\gamma};\leq_{\mathcal{S}^{\gamma}})$ for all $\gamma\geq\text{ch}(\Gamma)$. Our main result states that, if $\Gamma$ is finite, $\text{ch}(\Gamma)<\omega_{1}$ if $\Gamma\subset\omega$ and $\textnormal{ch}(\Gamma)<|\max\{\Gamma\}|^{+}$ otherwise.
@@ 2.1. Set theory
Therefore, for any ordinal of the form $\omega^{\sigma}$, we have a monoid structure $(\omega^{\sigma},\oplus,0)$.
Notation 2.4.
@@
By Lemma 2.3, $(\mathcal{S}^{\sigma};\oplus,0)$ is a submonoid of $(\omega^{\sigma};\oplus,0)$. For example, if $\Gamma=\{p_{1},\dots,p_{n}\}\subset\omega$, any $\alpha\in\mathcal{S}^{\sigma}$ can be written as
$\alpha=\omega^{\beta_{1}}\cdot(p_{1}\cdot a_{1}^{1}+\cdots+p_{n}\cdot a_{n}^{1})+\cdots+\omega^{\beta_{k}}\cdot(p_{1}\cdot a_{1}^{k}+\cdots+p_{n}\cdot a_{n}^{k}),$
for some ordinals $\beta_{k}<\cdots<\beta_{1}<\sigma$ and natural numbers $a_{i}^{j}$.
@@
For any ordinal $\alpha=\omega^{\alpha_{1}}\cdot a_{1}+\cdots+\omega^{\alpha_{n}}\cdot a_{n}$, we define the function $\overline{e}:\text{Ord}\to\text{Ord}$ by $\overline{e}(\alpha)=\alpha_{1}$. Given a set $\Gamma$ of ordinals, we define $\overline{e}(\Gamma)=\bigcup_{\gamma\in\Gamma}\overline{e}(\gamma)$.
@@
2.2. Game theory
Theorem 2.8 (Zermelo’s theorem).
Here, a finite game is a game in which any play of the game ends after a finite number of moves. In the case of games on partially ordered sets, we can ensure that the game is finite by requiring that there are no infinite antichains or infinite descending chains; that is, we shall work with well-partial orders.
@@ 2.2. Game theory
If $P$ is a well-partial order with a minimum and a maximum (different from each other), player $A$ has a winning strategy for Chomp on $P$.
@@
According to Zermelo’s theorem, one of the players must have a winning strategy. Suppose that this is player $B$. Then, after any first move of player $A$, player $B$ knows how to reply to win the game. However, if the first move of $A$ is to remove the maximum, any response from $B$ is a move that player $A$ could have made. That is, player $A$ could have used $B$’s strategy from the beginning. ∎
A reasoning like the one above is commonly referred to as a stealing argument. Another very common strategy in game theory (also in games without perfect information) is imitation. As an example, consider the following situation:
@@
Let $(\mathcal{T};\leq_{\mathcal{T}})$ be well-partial order. Then, player $B$ has a winning strategy for Chomp on the poset $\mathcal{T}^{2}=(\{0,1\}\times\mathcal{T})\cup\{0\}$ with $0\leq_{\mathcal{T}^{2}}(a,b)$ for all $(a,b)\in\{0,1\}\times\mathcal{T}$ and $(a,b)\leq_{\mathcal{T}^{2}}(c,d)$ iff $a=c$ and $b\leq_{\mathcal{T}}d$.
@@
Player $B$ simply has to copy player $A$’s moves but in the “opposite branch” of the poset. For example, if player $A$ chooses the element $(0,c)$, player $B$ will choose $(1,c)$ on the next move. In this way, the elements of the form $(a,b)$ will be exhausted just before one of player $A$’s turns, forcing player $A$ to choose 0, which is the minimum, and lose. ∎
The Sprague–Grundy theorem, which states that any impartial game of perfect information under the normal play convention (the last player able to move is the winner) is equivalent to playing a poset game where the poset is a well-order (this is the popular game Nim with one “heap”), formalizes the idea of the previous proof. Although this is a very useful tool, we will not use it in the rest of the paper.
3. Chomp on $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$
The goal of this section is to show the order-theoretic properties needed to see that the game Chomp on the posets $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$ can be interesting.
@@ 3. Chomp on $(\mathcal{S}^{\sigma};\leq_{\mathcal{S}^{\sigma}})$
Let $\sigma$ and $\alpha$ be ordinals such that $\sigma\geq 2$ and $\alpha<\sigma$. For every set $\{p_{1},\dots,p_{n}\}$ of natural numbers, the set
$\mathcal{S}^{\sigma,\alpha}:=\{\omega^{\alpha}\cdot(p_{1}\cdot a_{1}+\cdots+p_{n}\cdot a_{n})\mid a_{1},\dots,a_{n}\in\omega\}\subset\mathcal{S}^{\sigma},$
with the order $\leq_{\mathcal{S}^{\sigma}}\upharpoonright\mathcal{S}^{\sigma,\alpha}$ is isomorphic to $(\mathcal{S}^{1},\leq_{\mathcal{S}^{1}})$.
@@
The function $f:\mathcal{S}^{1}\to\mathcal{S}^{\sigma,\alpha}$ given by $f(n)=\omega^{\alpha}\cdot n$ is an isomorphism. ∎

### v3

# Hanf Numbers for Poset Games

@@ Abstract.
Given two well partial orders $(P;\leq_{P})$ and $(T;\leq_{T})$, each with a minimum element, we study the following question: which player has a winning strategy for Chomp on the poset $(P\times T;\leq_{P\times T})$? Here, $(P\times T;\leq_{P\times T})$ denotes the poset obtained as the Cartesian product of $P$ and $T$, equipped with the corresponding lexicographic order. The answer to this very natural question depends strongly on the specific choice of $(P;\leq_{P})$ and $(T;\leq_{T})$. For this reason, we restrict our attention to classes of posets given by powers of a fixed poset: $\{(P^{\sigma};\leq_{P^{\sigma}})\mid\sigma\in\mathrm{Ord}\}$. A fundamental fact about these classes of structures is that, if the second player does not have a winning strategy for all the posets in $\{(P^{\sigma};\leq_{P^{\sigma}})\mid\sigma\in\mathrm{Ord}\}$, there exists an ordinal $\xi$ such that the second player has a winning strategy on $(P^{\xi};\leq_{P^{\xi}})$ but not on $(P^{\gamma};\leq_{P^{\gamma}})$ for all $\gamma\geq\xi+1$. Determining the corresponding ordinal for this Hanf number-style property constitutes the main objective of this work.
Inspired by results of García-Marco and Knauer, we focus on classes of posets with a purely algebraic definition. These posets arise from submonoids (with respect to the natural sum, or Hessenberg sum) of ordinals of the form $\omega^{\sigma}$ and are generated by sets of ordinals. In the process, we provide a test to determine whether a finite set $\Gamma$ of ordinals indeed yields well partial orders, and, using set-theoretic techniques, we establish an upper bound for the ordinal $\xi$: if $\Gamma\subset\omega$, then $\xi<\omega_{1}$, and otherwise $\xi<|\bigcup\Gamma|^{+}$.
@@ 1. Introduction
In a poset game, two players take turns choosing elements from a partially ordered set and, on each turn, removing the chosen element and all those above it. The first player to run out of elements to choose loses. A particular case (and one of the most popular poset games) is the game Chomp: the poset is a rectangle of dimension $n\times m$ where each element is smaller than all the elements above it and to its right; the first player to eliminate the minimum (the element at the bottom left) loses. We can also define this game on any poset with a global minimum and, of course, by removing it, we obtain a poset game with the rules we explained at the beginning. The history of this combinatorial game dates back to 1952, when the Dutch mathematician Frederik Schuh introduced the “game of divisors”. Quite some time later, David Gale described in [Gal74] the Chomp game in the way just mentioned.
A natural question one may ask is the following: given two posets $(P;\leq_{P})$ and $(T;\leq_{T})$, each with a global minimum, what can we say about the Chomp game on the poset $(P\times T;\leq_{P\times T})$? Here, $P\times T$ denotes the set of pairs $(p,q)$ with $p\in P$ and $q\in T$, and $\leq_{P\times T}$ is the lexicographic order; that is, $(p,q)\leq_{P\times T}(p^{\prime},q^{\prime})$ iff $p\leq_{P}p^{\prime}$ or $p=p^{\prime}$ and $q\leq_{T}q^{\prime}$. More specifically, a similar question can be asked for successive powers $(P^{n+1};\leq_{P^{n+1}}):=(P\times P^{n};\leq_{P\times P^{n}})$ of $(P;\leq_{P})$, with $n\geq 1$.
@@
We will see that the answer to these questions depends heavily on the posets $(P;\leq_{P})$ and $(T;\leq_{T})$ that we choose. We will focus on addressing the second question and, in particular, we will work with a generalization of the Chomp game that was introduced in [GMK18]. There, the Chomp game is studied on the posets given by numerical semigroups $\mathcal{S}\subset\mathbb{N}$ with the order defined by $a\leq_{\mathcal{S}}b$ iff $b-a\in\mathcal{S}$. This is a huge generalization of the original game, so that, as in the classic Chomp, there is no known constructive way to find winning strategies, except in particular cases.
@@
Specifically, we propose a generalization of the game studied in [GMK18] to a set-theoretic context: as in $\mathbb{N}$, any additively indecomposable ordinal has a monoid structure on which we can define a partial order and, thus, play the Chomp game. In particular, we will focus on submonoids with natural sum (also called Hessenberg sum). The following sections will make clear that this family of submonoids is very interesting for our purposes. For example, we prove that it is possible to apply results from [GMK18] to this context to obtain constructive winning strategies. Apart from that, we will see that the algebraic definition of these posets allows us to use purely set-theoretic techniques to draw some conclusions about the game.
@@
Since this paper may attract readers interested in game theory or set theory, the following section presents some very basic facts from both disciplines that will be necessary later. In the third section, we present some basic results about Chomp on $(P\times T;\leq_{P\times T})$ and study some order-theoretic properties of the well partial order generators; that is, sets of ordinals that generate monoids in which we can define a natural well partial order structure. For instance, we provide a very general test to determine whether a finite set of ordinals is a well partial order generator. In the final section, we will investigate classes of well partial orders $\{(\mathcal{S}^{\tau};\leq_{\mathcal{S}^{\tau}})\mid\tau\geq\sigma\}$ generated by sets of ordinals $\Gamma$, where the sets $\mathcal{S}^{\tau}\subseteq\omega^{\tau}$ are monoids under natural sum. In particular, if the second player does not have a winning strategy on all the posets of the corresponding class, then there exists a successor ordinal $\xi+1$, which we denote by $\text{ch}(\Gamma)$, such that the second player has a winning strategy on $(\mathcal{S}^{\alpha};\leq_{\mathcal{S}^{\alpha}})$, for $\alpha\leq\xi$, but not on $(\mathcal{S}^{\gamma};\leq_{\mathcal{S}^{\gamma}})$ for all $\gamma\geq\xi+1$. Our main result states that, if $\Gamma$ is finite, $\text{ch}(\Gamma)<\omega_{1}$ if $\Gamma\subset\omega$ and $\textnormal{ch}(\Gamma)<|\bigcup\Gamma|^{+}$ otherwise.
@@ 2. Preliminaries
2.1. Game theory
Theorem 2.1 (Zermelo’s theorem).
Here, a finite game is a game in which any play of the game ends after a finite number of moves. In the case of games on partially ordered sets, we can ensure that the game is finite by requiring that there are no infinite antichains or infinite descending chains; that is, we shall work with well partial orders.
Proposition 2.2.
If $P$ is a well partial order with a global minimum and a global maximum (different from each other), player $A$ has a winning strategy for Chomp on $P$.
The proof of the preceding result uses what is known as strategy stealing: the idea is that, if player B had a winning strategy, then player A could remove the maximum element in the first move, and thereafter any response by player B is something player A could have played as well. More details can be found in [FR15]. The following is also a fundamental argument in poset games:
Proposition 2.3.
Let $(T;\leq_{T})$ be a well partial order. Then, player $B$ has a winning strategy for Chomp on the poset $T^{2}=(\{0,1\}\times T)\cup\{0\}$ with $0\leq_{T^{2}}(a,b)$ for all $(a,b)\in\{0,1\}\times T$ and $(a,b)\leq_{T^{2}}(c,d)$ iff $a=c$ and $b\leq_{T}d$.
Proof.
Player $B$ simply has to copy player $A$’s moves but in the “opposite branch” of the poset: if player $A$ chooses the element $(0,c)$, player $B$ will choose $(1,c)$ on the next move. In this way, the elements of the form $(a,b)$ will be exhausted just before one of player $A$’s turns, forcing player $A$ to choose 0 and lose. ∎
The Sprague–Grundy theorem, which states that any impartial game of perfect information under the normal play convention (the last player able to move is the winner) is equivalent to playing a poset game where the poset is a well order (this is the popular game Nim with one “heap”), formalizes the idea of the previous proof. Although this is a very useful tool, we will not explicitly use it in the rest of the paper.
@@ 2.2. Set theory
Therefore, for any ordinal of the form $\omega^{\sigma}$, we have a monoid structure $(\omega^{\sigma};\oplus,0)$.
Notation 2.7.
@@
By Lemma 2.6, $(\mathcal{S}^{\sigma};\oplus,0)$ is a submonoid of $(\omega^{\sigma};\oplus,0)$. For example, if $\Gamma=\{p_{1},\dots,p_{n}\}\subset\omega$, any $\beta\in\mathcal{S}^{\sigma}$ can be written as
$\beta=\omega^{\alpha_{1}}\cdot(p_{1}\cdot b_{1}^{1}+\cdots+p_{n}\cdot b_{n}^{1})+\cdots+\omega^{\alpha_{m}}\cdot(p_{1}\cdot b_{1}^{m}+\cdots+p_{n}\cdot b_{n}^{m}),$ (1)
for some ordinals $\sigma>\alpha_{1}>\cdots>\alpha_{m}$ and $b_{i}^{j}\in\omega$.
@@
Given an ordinal $\alpha=\omega^{\alpha_{1}}\cdot a_{1}+\cdots+\omega^{\alpha_{n}}\cdot a_{n}$, we define the function $\overline{e}:\text{Ord}\to\text{Ord}$ by $\overline{e}(\alpha)=\alpha_{1}$ and $\overline{e}(0)=0$. For a set $\Gamma$ of ordinals, we define $\overline{e}(\Gamma)=\bigcup_{\gamma\in\Gamma}\overline{e}(\gamma)$.
@@
3. Chomp on $(P\times T;\leq_{P\times T})$
3.1. Basic results
As mentioned in the introduction, we will work with posets that possess a global minimum and, specifically, with posets that correspond to the Cartesian product of two posets endowed with the lexicographic order. The first thing we can observe is the following simple result:
@@ 3.1. Basic results
Let $(P;\leq_{P})$ and $(T;\leq_{T})$ be well partial orders with global minima. If player $A$ has a winning strategy for Chomp on $T$, then player A also has a winning strategy for Chomp on $(P\times T;\leq_{P\times T})$.
@@
If 0 is the global minimum of $P$ and $t$ is a winning move for A when playing on $(T;\leq_{T})$, then $(0,t)$ is a winning move for A on $(P\times T;\leq_{P\times T})$: the remaining poset is isomorphic to the poset that results after the move $t$ in $(T;\leq_{T})$. ∎
Apart from the above, in general we cannot say much more:
Example 3.2.
Consider the following posets:
0
11
11
0
22
0
11
22
44
33
55
11
0
22
33
44
55
66
0
11
22
33
44
55
66
88
77
99
$P_{1}$
$P_{2}$
$P_{3}$
$P_{4}$
$P_{5}$
Figure 1. Hasse diagram of some posets.
  1. (1)
    If $P$ is $P_{1}$ and $T$ is $P_{2}$, then player A has the winning strategy on $(P;\leq_{P})$, player B has it on $(T;\leq_{T})$, and $(1,0)$ is a winning move for player A on $(P\times T;\leq_{P\times T})$.
  2. (2)
    If $P$ is $P_{4}$ and $T$ is $P_{2}$, then player A has the winning strategy on $(P;\leq_{P})$, player B has it on $(T;\leq_{T})$, and player B also has it on $(P\times T;\leq_{P\times T})$.
  3. (3)
    If $P$ is $P_{2}$ and $T$ is $P_{2}$, $P_{3}$ or $P_{5}$, then player B has the winning strategy on $(P;\leq_{P})$, $(T;\leq_{T})$, and $(P\times T;\leq_{P\times T})$.
  4. (4)
    If $P$ is $P_{3}$ and $T$ is $P_{2}$, then player B has the winning strategy on $(P;\leq_{P})$ and $(T;\leq_{T})$, but $(1,1)$ is a winning move for player A on $(P\times T;\leq_{P\times T})$. And the same holds if $P$ is $P_{3}$ and $T$ is $P_{3}$ or if $P$ is $P_{5}$ and $T$ is $P_{5}$: player A can win with the move $(1,3)$ in the first case and with $(1,1)$ in the second.
From now on we will work with “powers” of posets:
Definition 3.3.
Let $(P;\leq_{P})$ be a poset with global minimum 0. We define $(P^{\sigma};\leq_{P^{\sigma}})$ as follows:
  • If $\sigma=0$, $P^{0}=\{0\}$ and if $\sigma=1$, $(P^{1};\leq_{P^{1}})=(P;\leq_{P})$.
  • If $\sigma=\delta+1>1$, $(P^{\sigma};\leq_{P^{\sigma}})=(P\times P^{\delta};\leq_{P\times P^{\delta}})$.
  • If $\sigma$ is a limit ordinal, $P^{\sigma}=\bigcup_{\alpha\in\sigma}P^{\alpha+1}\backslash(\{0\}\times P^{\alpha})$. The order is defined as expected: for all $p,q\in P^{\sigma}$, if $p,q\in P^{\alpha}$ for some $\alpha\in\sigma$ and $p\leq_{P^{\alpha}}q$, then $p\leq_{P^{\sigma}}q$; if $p\in P^{\alpha}$ and $q\in P^{\beta}$ with $\alpha\in\beta\in\sigma$, then $p\leq_{P^{\sigma}}q$.
@@
If $(P;\leq_{P})$ is a well partial order with a global minimum 0, then, for any ordinal $\sigma$, $(P^{\sigma};\leq_{P^{\sigma}})$ is a well partial order and contains a global minimum.
@@
By transfinite induction. Cases $\sigma=0$ and $\sigma=1$ are trivial. Let us assume that $(P^{\sigma};\leq_{P^{\sigma}})$ contains no infinite antichains but $(P^{\sigma+1};\leq_{P^{\sigma+1}})$ contains an infinite antichain $A=\{(p_{i},q_{j})\mid i,j\in\omega\}$. Assume that there are infinitely many distinct $p_{i}$ in $A$, but then $\{p_{i}\mid i\in\omega\}\subset P$ would be an infinite antichain. Thus, there must be only finitely many $p_{i}$ in $A$. We may fix one of them, say $p_{k}$, such that $A_{k}=\{(p_{k},q_{j})\mid(p_{k},q_{j})\in A\}$ is an infinite antichain. However, $\{p_{k}\}\times P^{\sigma}$, with the order inherited from $(P^{\sigma+1};\leq_{P^{\sigma+1}})$, is isomorphic to $(P^{\sigma};\leq_{P^{\sigma}})$, so there cannot be infinitely many distinct $q_{j}$ either. That is, $A$ cannot be an infinite antichain. If $\sigma$ is a limit ordinal and for no $\alpha\in\sigma$ does $P^{\alpha}$ contain infinite antichains, then any infinite antichain $A\subseteq P^{\sigma}$ must contain elements $p$ and $q$ such that $p\in P^{\alpha}$ and $q\in P^{\beta}$ with $\alpha\in\beta\in\sigma$. But then $p$ and $q$ would be comparable. Similarly, we can show that $P^{\sigma}$ contains no infinite descending chains. Therefore, $P^{\sigma}$ is a well partial order. The global minimum can also be obtained by transfinite induction: in $P^{\sigma+1}$ the global minimum is $(0,\min(P^{\sigma}))$, and if $\sigma$ is a limit ordinal, it is 0 itself. ∎
From now on we may study the classes of structures $\{(P^{\sigma};\leq_{P^{\sigma}})\mid\sigma\in\mathrm{Ord}\}$ determined by a well partial order $(P;\leq_{P})$ with a global minimum. The following are fundamental observations:
@@
Let $(P^{\sigma};\leq_{P^{\sigma}})$ be a well partial order with a global minimum. If player $A$ has a winning strategy on $(P^{\alpha};\leq_{P^{\alpha}})$, then $A$ also has a winning strategy on every poset $(P^{\sigma};\leq_{P^{\sigma}})$ with $\sigma>\alpha$. Hence, if player $B$ has a winning strategy on $(P^{\sigma};\leq_{P^{\sigma}})$, player $B$ also has a winning strategy on $(P^{\alpha};\leq_{P^{\alpha}})$ for every $\alpha<\sigma$.
If $\delta$ is a limit ordinal and $B$ has a winning strategy on $(P^{\alpha};\leq_{P^{\alpha}})$ for unboundedly many $\alpha$ less than $\delta$, then $B$ has a winning strategy on $(P^{\delta};\leq_{P^{\delta}})$.
@@
The first statement is proved as in Proposition 3.1. The consequence is immediate and follows from Zermelo’s theorem.
For the second part, note that, then, $B$ has a winning strategy on $(P^{\alpha};\leq_{P^{\alpha}})$ for every $\alpha<\delta$. Player $A$’s first move must necessarily be some $p\in P^{\gamma}$ for some $\gamma<\delta$, so it will be a move in $P^{\gamma}$, where $B$ has a winning strategy. ∎
We can exploit this idea to work with families of classes of posets:
Proposition 3.6.
Let $S$ be a set of well partial orders with global minima. There exists an ordinal $\xi$ such that, for every $(P;\leq_{P})\in S$, player $B$ has a winning strategy on $(P^{\sigma};\leq_{P^{\sigma}})$ for every ordinal $\sigma$ iff $B$ has a winning strategy on $(P^{\xi};\leq_{P^{\xi}})$.
Proof.
Let us define
[…]
Take $\xi=\bigcup_{(P;\leq_{P})\in W}\alpha_{P}$, where $\alpha_{P}$ denotes the least ordinal $\gamma$ such that $A$ has a winning strategy on $(P^{\gamma};\leq_{P^{\gamma}})$. Clearly, if $B$ has a winning strategy on $(P^{\xi};\leq_{P^{\xi}})$, then $(P;\leq_{P})\not\in W$, i.e., $B$ has a winning strategy on $(P^{\sigma};\leq_{P^{\sigma}})$ for every $\sigma$. ∎
For example, $S$ could be the set of all such posets with cardinality less than or equal to a fixed cardinal $\kappa$.
The argument in the preceding proposition, as well as phenomena such as the one described in Lemma 3.5, are well known in model theory. In that context, the ordinal $\xi$ would correspond to a “Hanf number” for the property “$B$ has a winning strategy” (see [Bal09, Ch. 4]).
Notation 3.7.
Lemma 3.5 shows that if $B$ does not have a winning strategy on all the posets $(P^{\sigma};\leq_{P^{\sigma}})$, there must be an ordinal $\xi$ such that $B$ has a winning strategy on $(P^{\xi};\leq_{P^{\xi}})$ but not on $(P^{\gamma};\leq_{P^{\gamma}})$ for all $\gamma\geq\xi+1$. We let $\textnormal{ch}(P;\leq_{P})$ denote $\xi+1$ and, for a set $S$ of posets, $\text{ch}(S)=\bigcup_{(P;\leq_{P})\in S}\textnormal{ch}(P;\leq_{P})$.
3.2. Algebraic approach
We shall see in this section that the posets given by Lemma 2.10 fit perfectly into the classes of posets under consideration. We first require a simple preliminary lemma.